Tell whether the statement is true or false:
The data 6, 4, 3, 8, 9, 12, 13, 9 has mean 9.

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

Mean = Sum of all given observations/ number of observations
= (6 + 4 + 3 + 8 + 9 + 12 + 13 + 9)/8 = (64/8)
=8
Hence, the given statement is false.

​Thanks